Find the equivalent capacitance.

C1=5.0 µF, C2=3.5 µF, C3=8.0 µF, C4=1.5 µF, C5=0.75µF, C6=15 µF


1
Expert's answer
2021-01-06T13:57:11-0500

There are 2 types of placements of capacitors: in series and parallel.

The formulas for these:

i) In series "\frac{1}{C_s}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}+\frac{1}{C_5} +\frac{1}{C_6};"


ii) parallel "C_p=C_1+C_2+C_3+C_4+C_5+C_6;"


We examine both cases for this task:

i) "\frac{1}{C_s}=\frac{1}{5.0}+\frac{1}{3.5}+\frac{1}{8.0}+\frac{1}{1.5}+\frac{1}{0.75}+\frac{1}{15}" = "\frac{1}{C_s}=\frac{1}{5.0}+\frac{2}{7}+\frac{1}{8.0}+\frac{2}{3}+\frac{4}{3}+\frac{1}{15}" =


="\frac{168+240+105+1736}{840}=\frac{2249}{840}=2\frac{569}{840};"


ii) "C_p=5.0+3.5+8.0+1.5+0.75+15=33.75" µF